Daily Events on 2023/05/18 (Thu)

last updated at  2023/05/18 10:32:15 JST


<< prev day

@

next day >>


Station

Pass No.
20230518-

Contact time (UT)

Max. Elev.
 (deg)

Events

AOS

LOS

USC20

0200

09:01:25

09:14:19

50.1

OP/OG upload
XRT: TI-CMD
EIS: TI-CMD
DR playback (by OP)

SNT1

4010

09:52:21

10:04:30

20.9

OP/OG upload, XRT:TI-CMD, EIS:TI-CMD (backup)
AOCS: Upload pointing table

USC20

0300

10:40:11

10:49:54

11.4

DR playback (by OP)
XRT: Table Upload
EIS: Table Upload

USC20

0400

21:02:37

21:14:36

17.8

DR playback (by OP)

USC20

0500

22:39:54

22:52:20

32.9

DR playback (by OP)

Flare list(LM) : http://www.lmsal.com/solarsoft/last_events/


DR information

OP main: 2023/05/18 11:19:00 - 2023/05/20 11:24:00 (UT)

Remaining amount
(prediction)

2533 Mbits @ OP_START ( 2023/05/18 11:19:00 )


DR usage: The DR would become empty on May 19 at 5 UT if all telescopes took no images until that time.


SOT

XRT

EIS

DR allocation [%]

54

23

23

DR allocation [Mbits]

3848

1639

1639

DR TLMLMT [Mbits]

2256 @OP_START
-> 6011

2540

2540


daily note

EIS:
* If possible, move table upload to 3rd pass to prevent interruption of HOP 463 *
- HOP 463 (7:30-10UT until 22nd)
- Support HOP 412 (SOT latitudinal scan)
- Support HOP 393 (around 0UT on the 20th)
- AR obs

XRT:
- HOP 463
- HOP 412
- HOP 393
- HOP 349 & 448
- Stray light measurements
- AR obs

SOT:
- HOP 463
- HOP 412
- HOP 393
- AR obs

POINTING:
2023/05/18 12:12:00.000  |    -18  |  -1002  |  HOP412  01/20
2023/05/18 12:35:00.000  |    -18  |   -902  |  HOP412  02/20
2023/05/18 13:50:30.000  |    -18  |   -802  |  HOP412  03/20
2023/05/18 14:13:30.000  |    -18  |   -702  |  HOP412  04/20
2023/05/18 15:28:30.000  |    -18  |   -602  |  HOP412  05/20
2023/05/18 15:51:30.000  |    -18  |   -502  |  HOP412  06/20
2023/05/18 17:16:00.000  |    -18  |   -402  |  HOP412  07/20
2023/05/18 17:39:00.000  |    -18  |   -302  |  HOP412  08/20
2023/05/18 18:52:30.000  |    -18  |   -202  |  HOP412  09/20
2023/05/18 19:15:30.000  |    -18  |   -102  |  HOP412  10/20
2023/05/18 20:29:00.000  |    -18  |    -12  |  HOP412  11/20
2023/05/18 20:52:00.000  |    -18  |     88  |  HOP412  12/20
2023/05/18 22:06:30.000  |    -18  |    188  |  HOP412  13/20
2023/05/18 22:29:30.000  |    -18  |    288  |  HOP412  14/20
2023/05/18 23:42:00.000  |    -18  |    388  |  HOP412  15/20
2023/05/19 00:05:00.000  |    -18  |    488  |  HOP412  16/20
2023/05/19 01:18:30.000  |    -18  |    588  |  HOP412  17/20
2023/05/19 01:41:30.000  |    -18  |    688  |  HOP412  18/20
2023/05/19 02:57:00.000  |    -18  |    788  |  HOP412  19/20
2023/05/19 03:20:00.000  |    -18  |    888  |  HOP412  20/20
2023/05/19 03:43:00.000  |      0  |      0  |  End of HOP412
EOP


[Maintenance Note (Yamagishi), for CO/CP/RT Tohban]
Some software packages will be updated and security patch will be installed on machines
'eisco', 'sotcows1', 'sotcows2', 'aocswin', 'higashirinkan', and 'ope-gw'
at 15:00-17:30(JST) on 18th May. When that's over, these machines will be rebooted.
Please do not use these machines during updating, and report if you have any trouble.


CHIEF
TOHBAN

SUB
TOHBAN

CP

SOT CO

XRT CO

EIS CO

EVENING

Kotanagi

Kawabata

Kawabata->Shine (RCO)

Fleming(RCO)

Brooks

MORNING


operation files

pointing file : cmdpln/20230518/re-point_202305180200.txt